### v1

# LLaMIPa: An Incremental Discourse Parser

Abstract
This paper provides the first discourse parsing experiments with a large language model (LLM) finetuned on corpora annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). The result is a discourse parser, LLaMIPa (LLaMA Incremental Parser), which is able to more fully exploit discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units. Furthermore, it is able to process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
1 Introduction
Understanding a conversation requires more than just grasping the contents of its utterances. One must also understand how different utterances relate to one another: whether an utterance answers a question or serves to explain or correct something else that was said, for example. Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges or separate nodes.
Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inference required to predict the graph elements.
Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2018), which offer richer representations of linguistic content, and can address factor (ii). But given their limited input size, they are confined to look at one small part of a discourse at a time and so cannot really address factor (iii).
Generative large language models (LLMs) can substantially improve discourse parsing. They not only encode large amounts of lexical and conceptual information, addressing problem (ii) but their increased context windows permits an incremental parsing strategy to address problem (iii). Such a strategy makes sense cognitively and linguistically, as prior discourse structure clearly has an effect on inferences about which subsequent discourse units are attached (the link task) and how they are attached—i.e. what semantic relations link them (the link+relation task) Asher and Lascarides, (2003). Further, an incremental approach naturally applies to downstream applications like conversational agent, in which the information about the previously inferred discourse structure must be updated and available in real time.
We develop an incremental parsing model LLaMIPa (LLaMa Incremental Parser), by fine-tuning LLaMA3, that uses previously inferred discourse structure to simultaneously predict the link and the link+relation tasks between new discourse units. We show LLaMIPa substantially outperforms the state of the art on both tasks.
We demonstrate LLaMIPa’s performance using three datasets annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multi-party chats from an online game Asher et al., (2016). We further test (STAC-based) LLaMIPa on an out of domain dataset, Molweni Li et al., (2020).
Not only does LLaMIPa outperform prior models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i). In particular, LLaMIPa can identify cases in which a single EDU has two parent EDUs—a situation that arises naturally in both multiparty and situated dialogue where a speaker can respond to multiple channels (conversational or multimodal) at once. Most current SDRT based parsers cannot handle such structures Shi and Huang, (2019); Liu and Chen, (2021); Wang et al., (2021); Chi and Rudnicky, (2022).
In what follows, Section 2 provides background on discourse structure. Section 3 presents related work, while Section 4 details the multiparty corpora we use for training and testing. Sections 5 and 6 describe our model and results. Section 7 describes ablation experiments, and then we conclude in Section 8.
2 Background
Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999). Such relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in the discourse graph.
There are two main theories that have investigated complete discourse structures for texts: Rhetorical Structure Theory (RST; Mann and Thompson,, 1987) and SDRT (Asher,, 1993; Asher and Lascarides,, 2003). RST builds discourse trees over EDUs with the semantic relation labels on the head nodes dominating subtrees.
SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) structure in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs). EEUs represent salient aspects of the nonlinguistic context in which a conversation takes place, and may be connected to the discourse DAG. SDRT also allows subgraphs, or complex discourse units (CDUs) to act as terms of semantic relations.
Discourse parsing requires addressing three tasks. The segmentation task involves decomposing a text and (descriptions of) actions into EDUs and EEUs. The link task involves identifying pairs of EDUs and/or EEUs that are semantically related via discourse relations—that is, the arcs in the discourse graph. The link+relation task involves identifying the labels for each attached pair.
3 Related work
Traditionally with SDRT style parsers, gold segmentation is assumed Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked and then labels are subsequently predicted for these links. While we assume gold segmentation, our proposed architecture combines the attachment and labeling tasks.
Several strategies have evolved to take account of the complex information required for discourse parsing within a dependency based paradigm like that proposed by SDRT or DLTag (Forbes et al.,, 2003). Research has shown that semantic features of EDUs are crucial for solving the link+relation task Mann and Thompson, (1987); Asher and Lascarides, (2003). Older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) thus used manual features about EDU pairs and simple ML models to build a local model to predict link and link+relation label in a pipeline fashion. A local model uses information just about pairs of EDUs that are to be attached; it does not exploit information about the discourse structure around that pair. Afantenos et al., (2015); Perret et al., (2016) also added decoding mechanisms with post hoc constraints on the output structures, following Muller et al., (2012). Although exploiting little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023). We do a detailed comparison between an encoder based local parsing model and LLaMIPa.
Shi and Huang, (2019) were the first to obtain significant discourse parsing results using a deep learning architecture on STAC. They attempted to capture incremental and contextual effects in their model by training a supplemental Structured Encoder to incrementally update attachment paths (sequences of parent-child EDUs). However, Wang et al., (2021) showed that the model obtains similar scores with or without this encoder.
Liu and Chen, (2021) develop an approach that uses a pre-trained RoBERTa model Liu et al., (2019) to provide embeddings for EDUS. In addition, they use a bi-GRU cell Cho et al., (2014) to capture contextual information; but this requires them to limit the size of the input. Their model uses two linear layers for link and relation prediction.
Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and edge embeddings. The model uses a Hierachichal GRU gate for contextual EDU representations. They then apply the SSA-GNN to capture implicit structural information between EDUs, using a Structure-Aware Scaled Dot-Product Attention Zhu et al., (2019); Wang et al., (2020) to update edge and EDU representations. They further train a teacher network to supplement the standard classification loss with an auxiliary loss. Chi and Rudnicky, (2022) developed a more complex model based on Shi and Huang, (2019).
None of the models by Shi and Huang, (2019); Liu and Chen, (2021); Guz et al., (2020); Wang et al., (2021); Chi and Rudnicky, (2022) predict multiple parents for an EDU (or EEU). The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023); Thompson et al., (2024)’s model BERTLine. BERTLine uses BERT embeddings for EDUs and EEUs, along with a linear layer for features like speaker change and distance information, to build a local model and predicts link and link+relation in pipeline fashion. However the relation labeling uses a multitask architecture to capture informational dependencies between link and relation labeling decisions. Bennis et al., (2023) show that their model is competitive with approaches that try to build in more contextual knowledge Liu and Chen, (2021); Wang et al., (2021). Thompson et al., (2024) uses BERTLine on the MSDC, which makes it a natural baseline for our work on situated, multiparty conversation.
Recent work has investigated using LLMs to predict discourse structure, by finetuning LLaMA2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024). The reported results on RST corpora are not directly comparable to ours, but they are significantly lower numerically. Chan et al., (2024); Yung et al., (2024); Fan et al., (2024); Henning and Stede, (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations. The results are significantly worse than dedicated discourse parsers; testing with 3-shot learning on a version of the STAC corpus (STAC-L; see Section 4), Chan et al., (2024) reports an F1 score of 21.3 for link prediction and 7.4 for relation prediction by ChatGPT.
4 Data sets
MSDC. The MSDC provides full discourse structures in the style of SDRT to dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019). In the MDC, dialogues involve a builder, whose task is to build block formations in a simplified Minecraft world, and an architect, who guides the builder by giving instructions through chat. The MSDC annotates discourse relations between both chat moves (EDUs) and linguistic descriptions of the builder’s pick and place actions (EEUs), provided by the MDC logs.
The MSDC’s multimodal nature presents many opportunities to simultaneously respond to a conversational move and a building move, and indeed, the MSDC has numerous discourse units with multiple parents (MPDUs). Figure 1 shows an excerpt from the MSDC with two examples. For instance, after placing two blocks, the builder asks “like that?”. The architect’s reply, “not quite,” not only serves to answer the builder’s question but simultaneously to correct the builder’s last moves. The following two elaborating discourse moves fill out the full content of the Correction.
Figure 1 shows another important aspect of the MSDC, namely that builders tend to perform sequences of pick and place moves uninterrupted by chat moves. These action sequences are grouped together as CDUs in the MSDC, indicated by dashed boxes in Figure 1. Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent them from drowning out the more complex and interesting linguistic interactions in the corpus. We used the MSDC with this modification.
| DU-type | STAC-L | STAC-Sit | MSDC |
| Train/Test | Train/Test | Train/Test | |
| EDU | 11398/1154 | 11707/1163 | 17135/5402 |
| EEU | 0/0 | 11566/1156 | 4687/1473 |
| MDPU | 838//86 | 999/96 | 4789/1476 |
| MPDU-3 | 94/9 | 95/6 | 85/35 |
| MPDU>3 | 11/0 | 7/0 | 1/3 |
STAC. STAC111https://https://www.irit.fr/STAC/corpus.html is a corpus of multiparty chats from the game Settlers of Catan, in which players trade or otherwise acquire resources in order to build roads and settlements and thereby score victory points. STAC has two versions: STAC-L, in which only relations between chat moves (EDUs) are annotated in the absence of the context provided by nonlinguistic events, and a situated version, STAC-Sit, in which both chat moves and (descriptions of) game moves (EEUs) are annotated in the presence of an integrated linguistic and nonlinguistic context.
STAC contains MPDUs due to both multimodal interactions and multiparty discourse, in which one player can simultneously respond to multiple players at once. Relevant statistics for STAC and the MSDC are found in Table 1.
STAC-Sit, like the MSDC, also contains sequences of game-only moves. Bennis et al., (2023) compressed such sequences into individual EEUs and we follow this choice. Unlike the MSDC, however, STAC contains numerous CDUs for chat moves and for chat/game interactions (as opposed to CDUs of action sequences only). Following Bennis et al., (2023) and most of the literature on SDRT parsing, we did not treat CDUs directly but used the “flattening" procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs. A final simplication of the original STAC corpus involved removing EEUs that neither interact with EDUs n or figure in EEU sequences that interact with EDUs. This choice, like the one to compress EEU sequences, was made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
Molweni. The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015) a collection of exchanges from an online forum for discussing issues related to Ubuntu. Unlike the MSDC and STAC, Molweni does not contain CDUs, EEUs, or MPDUs. We used the version of Molweni Test re-annotated by Li et al., (2024).
5 LLaMIPa: An Incremental Parser
LLaMIPa leverages the sequence to sequence architecture and the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy. As it generates a discourse graph for a conversation , the LLM should base its link and relation predictions for an EDU in on the graph it has already constructed for EDUs preceding .
Formally, let be a conversation consisting of a sequence of EDUs or EEUs. Our LLM takes as input a set of of discourse and (possibly) event units and a discourse graph . At step 0, is , the head EDU, and , the null discourse graph. After that, is a set , where 1 and where is the window size, set to ; and contains links for . At each step, the model predicts a graph that extends by integrating link+relation predictions for all of the EDUs in the next dialogue turn or, if the next move is an action sequence, the EEU representing this sequence. This process repeats until LLaMIPa arrives at the final unit of , effectively creating a “sliding context window” that will eventually slide over an entire dialogue, permitting the model to potentially exploit a large amount of contextual information.
We represent relation instances as EDU index pairs connected by a relation type, e.g. RES(0,1), for the instance of a Result relation holding between EDUs 0 and 1. During training, each subsequent instance in a dialogue adds to the context the previous turn EDUs and EEUs, but we take the gold, rather than predicted, discourse structure between those units to make our input graph . is specified as a list of the relations between these units in the same format as the predicted relations.
We used a sliding context widow of 15 EDUs to optimize the number of training samples while capturing the most number of discourse links that have statistical support. As discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than 15 in the SDRT corpora.
Using our corpora formatted for incremental relation prediction, we finetuned LLaMA-2 Touvron et al., (2023), and LLaMA-3222https://llama.meta.com/llama3/ to produce LLaMIPa2 and LLaMIPa3. In order to do this without training the full model weights, we use QLoRA Dettmers et al., (2024) to train for 3 epochs. The model parameter details are in Table 6 in Appendix A. The finetuned models will be made publicly available.
Once LLaMIPa is finetuned on the training corpus, the generation procedure uses predicted discourse structure instead of the gold structures during inference (See Figure 2). Proceeding turn by turn, LLaMIPa produces a complete discourse structure for a dialogue in which both links and relations are predicted. The LLaMIPa3 model and generation code is available here333https://huggingface.co/Kqte/LLaMIPa.
6 Results
We first ran our finetuning experiments on the MSDC corpus, and assessed LLaMIPa’s performance on link and relation type prediction when we include the gold dialogue structure in each sample context. We then tested LLaMIPa by using the previously predicted relations in the context instead of the gold ones, as in the generation step described above, in order to have our incremental parser build a predicted structure from beginning to end. Table 2 shows that LLaMIPa3 significantly outperforms BERTLine and outperforms LLaMIPa2 both using gold structure (LLaMIPa3+g vs. LLaMIPa2+g) and predicted structure (LLaMIPa3+p vs. LLaMIPa2+p). Below we use the better performing LLaMIPa3+p for other evaluations.
To test the generalizability of the model, we also finetuned LLaMIPa3 on STAC-Sit and compared it to BERTLine on MSDC and STAC-Sit. For both link prediction and link relation prediction, LLaMIPa shows a gain of almost 10 F1 points or 11% on link and 10.5 F1 points or 15% on link relation labeling. LLaMIPa posted even more impressive gains for predictions of certain relations. When we look at F1 breakdown by relation type in Table 5, the largest relative increase comes from improved predictions for the Narration and Correction relation types, which Thompson et al., (2024) cite as the most difficult to predict given their particular characteristics.
We then compared our model to a broader range of SDRT discourse parsers on the STAC linguistic version (STAC-L), as previous work focused on this version of the corpus. Table 4 shows that LLaMIPa3 finetuned on STAC-L and evaluated on STAC-L Test scored significantly higher on link prediction and got a statistically significantly higher score on link + relation prediction over the state-of-the-art parser Chi and Rudnicky, (2022).
Nevertheless, LLaMIPa3 has substantially lower performance on STAC-L (linguistic) than on either of the situated corpora. This mirrors human efforts at annotation. Asher et al., (2016, 2021) report that annotating the STAC linguistic data set was more difficult; annotators felt that information needed to determine the discourse structure was missing.
We also investigated LLaMIPa3 for transfer learning, finetuning on STAC-L and testing on the Molweni Cleaned Test corpus of Li et al., (2024). While superior to other models, in comparison to Li et al., (2024), LLaMIPa3 fared worse on link prediction but was far superior with respect to link + relation prediction.
Overall, LLaMIPa has several advantages. It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously. Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023). Our model also outperforms prior models on all relevant corpora and pretty much all tasks we studied. We hypothesize that this performance is due to the richer contextual information supplied to LLaMIPa, a hypothesis we now explore further.
| MSDC | STAC Situated | |||
| Link | Link+Rel | Link | Link+Rel | |
| LLaMIPa2+g | 0.8561 | 0.7664 | - | - |
| LLaMIPa2+p | 0.8579 | 0.7570 | - | - |
| LLaMIPa3+g | 0.9004 | 0.8154 | - | - |
| LLaMIPa3+p | 0.8830 | 0.7951 | 0.8612 | 0.7796 |
| BERTLine | 0.7870 | 0.6901 | 0.7667 | 0.6788 |
| Distance | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 |
| LLaMIPa-3 | 0.90 | 0.88 | 0.82 | 0.76 | 0.82 | 0.80 | 0.79 | 0.71 | 0.90 | 0.76 | 0.50 | 0.40 | 1.0 | n/a |
| BERTLine | 0.70 | 0.61 | 0.43 | 0.32 | 0.38 | 0.02 | 0.21 | 0.28 | 0.08 | n/a | n/a | n/a | n/a | n/a |
| STAC-L | STAC- L/Molweni-clean | |||
| Link | Link+Rel | Link | Link+Rel | |
| Perret et al., (2016) | 0.69 | 0.531 | - | - |
| Shi and Huang, (2019) | 0.719 | 0.536 | - | - |
| Liu and Chen, (2021) | 0.731 | 0.571 | - | - |
| Wang et al., (2021) | 0.734 | 0.573 | - | - |
| Fan et al., (2022) | 0.736 | 0.574 | - | - |
| Chi and Rudnicky, (2022) | 0.744 | 0.596 | 0.645 | 0.380 |
| Li et al., (2024) | 0.593 | 0.386 | 0.756 | 0.312 |
| Bennis et al., (2023) (BERTLine) | 0.730 | 0.562 | - | - |
| Fan et al., (2024) (ChatGPT) | 0.599 | 0.252 | - | - |
| Chan et al., (2024) (ChatGPT) | 0.213 | 0.074 | - | - |
| LLaMIPa-3+p | 0.775 | 0.607 | 0.712 | 0.405 |
| BERTLine | LLaMA3-Local | LLaMIPa3+p | LLaMIPa3+g | LLaMIPa-rand | No Structure | |
| Result | 0.85 | 0.86 | 0.90 | 0.91 | 0.83 | 0.72 |
| Acknowledgement | 0.81 | 0.85 | 0.86 | 0.86 | 0.81 | 0.72 |
| Narration | 0.50 [0.73] | 0.54 | 0.82 | 0.91 | 0.28 | 0.18 |
| Elaboration | 0.75 | 0.77 | 0.77 | 0.77 | 0.75 | 0.68 |
| Correction | 0.31 | 0.64 | 0.73 | 0.80 | 0.49 | 0.52 |
| Continuation | 0.44 | 0.47 | 0.49 | 0.50 | 0.44 | 0.33 |
| Question-answer Pair | 0.76 | 0.80 | 0.82 | 0.83 | 0.42 | 0.14 |
| Comment | 0.50 | 0.57 | 0.60 | 0.61 | 0.57 | 0.54 |
| Confirmation-Question | 0.86 | 0.89 | 0.93 | 0.93 | 0.91 | 0.89 |
| Clarification-Question | 0.61 | 0.66 | 0.73 | 0.73 | 0.70 | 0.41 |
| Contrast | 0.80 | 0.79 | 0.75 | 0.74 | 0.72 | 0.68 |
| Question-Elaboration | 0.39 | 0.36 | 0.30 | 0.36 | 0.33 | 0.33 |
| Alternation | 0.88 | 0.88 | 0.83 | 0.88 | 0.90 | 0.96 |
| Explanation | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Conditional | 0.58 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Sequence | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 |
| Link+Rel F1 | 0.69 [0.71] | 0.73 | 0.80 | 0.81 | 0.65 | 0.56 |
| Link F1 | 0.78 [0.80] | 0.82 | 0.88 | 0.90 | 0.77 | 0.72 |
7 The role of context
A particularly remarkable demonstration of LLaMIPa’s discourse parsing prowess is its success in predicting Narration instances in the MSDC. Narrations are frequently quite long (with a significant proportion above distance 10), and so difficult to capture by local parsers like BERTLine Thompson et al., (2024). Table 3 shows Narration F1 at distances between 3 and 14 EDUs. Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance 10. As explained in Section 5, LLaMIPa allows for training up to distance 15, and impressively, it continues to make (somewhat steadily good) predictions up to distance 15.
LLaMIPa also has very strong performance in predicting instances of the Correction relation, which also has a significant number of long distance instances in the MSDC. Table 5 shows that its performance is more than double BERTLine’s on this relation; its F1 score for Correction is over 42 F1 points more than BERTLine’s.
We hypothesize that LLaMIPa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context. It can use the meaning of previous discourse moves as well as an explicit representation of their discourse structure. This hypothesis coheres with our understanding of how Narration functions in the MSDC. Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1). While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements. This pattern is regular enough that one can find some Narrations algorithmically as in Thompson et al., (2024)’s “Second Pass”: taking predictions from a first pass of the local model, Thompson et al., then apply an algorithmic approach to predict long distance Narration links from the local model’s predictions. We also tried this second pass for LLaMIPa, however there was no improvement on F1 for Narration scores, which shows that LLaMIPa had already learned this rule during training.
Besides its performance on Narration and Correction, LLaMIPa offers significant improvements over BERTLine for many other relation types, even those whose instances are typically not very long distance. Table 5 shows LLaMIPa has significantly higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question and Comment. We hypothesize that predictions on these too profit from complex contextual information.
LLaMIPa produces poorer results for relations such as Contrast, Alternation and Question-Elaboration. This may be because instances of these relations can be inferred reliably just from information given by the related EDUs or EEUs.
First ablation: embeddings vs. context. While the previous discussion suggests that LLaMA’s larger context window is responsible for its better performance, we wanted to check this hypothesis further. So we finetuned LLaMA3-8b over pairs of EDUs as input to predict links and relation types between pairs of EDUs. This produces a local model roughly comparable to BERTLine’s, though BERTLine’s architecture is different. Table 5 shows that, compared to BERTLine, LLaMA3-Local gave stronger performance without using the prior discourse context for link and relation prediction. However, when compared to LLaMIPa3, LLaMA3-Local has significantly lower link and link+relation labeling predictions overall. LLaMIPa3’s relation predictions for many relations are also significantly better—indeed sometimes vastly superior as in the F1 for Narration. This supports our conclusion that although LLaMA’s superior embeddings play a role, many relations, and especially relations like Narration, require more contextual information like that available to LLaMIPa.
Further ablations: removing structure. Given the result of our first ablation experiment, we wanted to investigate further whether LLaMIPa was using the discourse structure for its predictions, or was solely relying on the longer input sequence of EDUs and EEUs.
Our set up for LLaMIPa makes the assumption that it is the prior discourse structure that is relevant. In Table 5, there is a small but consistent drop between gold and predicted contexts in F1 performance for all relation types. This supports the conclusion that LLaMIPa takes the discourse structure of its context into account and that the quality of that structure affects prediction.
To provide more support for this conclusion, we performed another ablation in which we perturbed the structure at each increment by randomizing both the relation type and EDU index for each relation. We then tested LLaMIPa (finetuned in the usual way) on this input. LLaMIPa-rand showed a significant drop in performance from LLaMA3-Local for link and across almost all relation types (with the exception of Contrast and Question Elaboration). The drop for Narration and Correct is sharp, as well as for Question-answer Pair. We then performed a variant ablation, in which we removed all structure from LLaMIPa’s input context, leaving only a sequence of EDUs and EEUs. The performance of LLaMIPa on this data set (in the column “No Structure” of Table 5) shows another significant drop both in link and in link+relation predictions. The drop in relation labeling in addition holds for almost all relations, and a dramatic drop, especially for Narration, Correction, Clarification Question and Question-Answer Pair. This provides strong evidence that LLaMIPa needs to the use discourse structure and not just the EDU/EEU input sequence of the context to optimize its predictions for links and relation labels and that discourse structure is crucial for predicting some relations.
8 Conclusions
In this paper, we presented LLaMIPa3, an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for new discourse units. We finetuned both LLaMa2-13b and LLaMA3-8b models to produce LLaMIPa2 and LLaMIPa3. We showed that LLaMIPa3 performs better than other parsers that use SDRT data sets. It excels in particular on data sets that feature conversation that contains both linguistic moves and nonlinguistic contexts that help interpret them.
We also compared LLaMIPa3 with a local model discourse parser based on LLaMA3-8b as well as on variants of LLaMIPa3 that had differing levels of access to the discourse context. LLaMIPa3 had significantly higher F1 scores for attachment and relation labeling on the data sets we used for evaluation than all of these.
The differential performance of LLaMIPa3 on STAC-L versus STAC-Sit, which reflects impressions of human annotators, suggests that nonlinguistic contexts are important for stable interpretation of situated conversation. We humans typically invent a nonlinguistic context for a conversation when one isn’t given. But such an invention is subjective and may contribute to variations in understanding and prediction. We suspect something like this might be true for LLMs too. In future work, we hope to use LLaMIPa3 to help conversational robots to improve their performance in interactive tasks.
Limitations
We did not exhaustively test all models on the situated data sets. Given the number of MPDUs especially in the MSDC, however, we felt that their performance would not compare to LLaMIPa’s. We plan to do more extensive testing in the future to compare models.
LLaMIPa, like previous parsers, presumes that the edus/eeus are already segmented. However extant work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
Though ablations show that LLaMIPa uses discourse structure, it remains unclear how it leverages that structure to predict relations like Narration, Corrections, and QAP.
Like all supervised models, LLaMIPa is limited to making predictions on data sets that are similar to its training data. Given the heavy annotation effort required to build data sets annotated for discourse structure, this arguably puts a bottleneck on the generalizability of our approach to data sets in languages other than English and for conversation styles, such as multiparty spoken conversation, that are significantly different from the chat-based dialogues in the MSDC and STAC. Future work will explore efficient ways of harnessing extant data sources through, for instance, automatic translation, data augmentation and or/heuristics, to facilitate the transfer to other languages and domains.
Ethics Statement
Our work here has been to improve the capacities of AI models to understand the semantic structure of discourse and conversation. We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable models to eventually produce less pragmatically grounded or conversational hallucinations.
Appendix A Appendix: Model Parameters
| GPUs | |
| 4 NVIDIA Volta V100 | |
| Hyperparameters | |
| Training epochs | 3 |
| batch size | 4 |
| optimizer | Adam |
| learning rate | 2e-4 |
| learning rate scheduler | linear warm-up and |
| cosine annealing | |
| warm-up ratio | 0.03 |
| gradient clipping | 0.3 |
| lora r | 64 |
| lora (alpha) | 16 |
| lora dropout ratio | 0.1 |
| lora target modules | Only Attention Blocks |
| (q_proj, v_proj) | |
| quantization for LLaMA3 | 4-bit NormalFloat |
Table 6 gives the hyperparameters used for finetuning LLaMIPa-2 and LLaMIPa-3 along with the computing resources. We adapted the finetuning code from the following repository444https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb. We provide LLaMIPa3 trained on MSDC as well as the MSDC test data here555https://huggingface.co/Kqte/LLaMIPa
Appendix B Appendix: STAC Linguistic data
We constructed STAC-L, our STAC linguistic corpus, from the STAC data available on the STAC website (https://https://www.irit.fr/STAC/corpus.html) and adopted standard procedures for flattening CDUs. However, we found that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs. Shi and Huang, (2019) also report a different number of EDUs from what we have and from what Bennis et al., (2023) report. When we finetuned and tested LLaMIPa on the STAC linguistic only data available in the BERTLine repository 666https://github.com/zineb198/LineBert, we had lower scores than what we report in Table 4. LLaMIPa-3+p had an F1 for link of 75.4 and a link+relation F1 of 57.7 on this version of STAC linguistic. On the transfer learning from STAC linguistic to Molweni Cleaned Test, LLaMIPa-3+p had an F1 of 68.5 for link and 38.7 for link+relation.